# Condor: A Code Discriminator Integrating General Semantics with Code Details

### v1

@@ Abstract.
LLMs demonstrate significant potential across various software engineering tasks. However, they still face challenges in generating correct code on the first attempt when addressing complex requirements. Introducing a discriminator to select reliable outputs from multiple generated results is an effective way to enhance their reliability and stability. Currently, these discriminators fall into two categories: execution-based discriminators and non-execution-based discriminators. Execution-based discriminators face flexibility challenges due to difficulties in obtaining test cases and security concerns, while non-execution-based discriminators, although more flexible, struggle to capture subtle differences in code details. To maintain flexibility while improving the model’s ability to capture fine-grained code details, this paper proposes Condor. We first design contrastive learning to optimize the code representations of the base model, enabling it to reflect differences in code details. Then, we leverage intermediate data from the code modification process to further enrich the discriminator’s training data, enhancing its ability to discern code details. Experimental results indicate that on the subtle code difference dataset (i.e., CodeNanoFix), Condor significantly outperforms other discriminators in discriminative performance: Condor (1.3B) improves the discriminative F1 score of DeepSeek-Coder (1.3B) from 67% to 73%. In discriminating LLM-generated outputs, Condor (1.3B) and Condor (110M) raise the Pass@1 score of Meta-Llama-3.1-Instruct (70B) on the CodeNanoFix dataset from 52.64% to 62.63% and 59.64%, respectively. Moreover, Condor demonstrates strong generalization capabilities on the MBPP and APPS datasets. For example, Condor (1.3B) improves the Pass@1 of Meta-Llama-3.1-Instruct (70B) on the APPS dataset by 147.05%.
Code Discriminator, Code Embedding, Code Generation, Code Repair
copyright: noneccs: Software and its engineering Automatic programming
1. Introduction
@@
With the widespread application of large language models (LLMs), their potential for tasks in software engineering (e.g., code generation and code repair) receives significant attention (Guo et al., 2024; Dubey et al., 2024; Team, 2024; Nam et al., 2024; OpenAI, 2024). However, they typically struggle to produce correct code on the first attempt when addressing complex requirements such as advanced algorithmic problems (Hendrycks et al., 2021; Yetiştiren et al., 2023; Khan et al., 2023; Ji et al., 2023). Existing approaches typically employ an automated discriminator to select a correct result from multiple outputs generated by LLMs, improving the reliability and stability of LLMs when handling complex problems (Inala et al., 2022).
@@
Currently, these approaches for code discrimination fall into two main categories: execution-based approaches and non-execution-based approaches. Execution-based approaches require running the code before making a discriminative decision, utilizing feedback from execution as additional detailed information to enhance the discriminator’s performance. However, execution-based models face challenges in flexibility. First, executing code poses challenges because (1) many real-world downstream reasoning tasks lack test cases (Briand, 2007), (2) LLM-generated code may contain incomplete snippets (Chen et al., 2021a), and (3) the testing environments and other dependencies are not always accessible (Inala et al., 2022; cop, 2024). Second, even when execution is possible, running LLM-generated code raises safety concerns. On the other hand, non-execution-based approaches do not execute code and rely solely on the details presented in the code text, providing greater flexibility and applicability across different application scenarios.
@@
However, existing non-execution-based discriminators typically leverage general code representation capabilities without adapting to the nuanced details of the programming process (Inala et al., 2022; Zhang et al., 2023). Challenges remain in improving the discriminator’s effectiveness at capturing intricate code details. In the development scenario, the functionality of code often depends on key minor changes, such as variable updates, conditional logic modifications, or adjustments in function calls. For example, as shown in Figure 1, the upper section illustrates a problem description about finding a number K equally distant from two distinct numbers A and B. Submission 1 displays buggy code, while submission 2 shows the correct code after the user’s modification. Notably, only three characters differ between the two submissions (i.e., ‘(’, ‘)’, and ‘/’), yet the functionality of the first submission remains incorrect. Similarly, outputs generated by LLM can also contain errors caused by subtle differences, demonstrating that such mistakes are not unique to humans. For example, as shown below Figure 1, the model (i.e., Meta-LLama-3.1-Instruct-70B) generates two outputs for Task 173 in the MBPP (Austin et al., 2021) benchmark, one of which is incorrect due to a minor substitution, where ‘all’ is mistakenly replaced with ‘any’. Existing discriminators primarily rely on embeddings provided by other general-purpose code models as a basis for representation. This limits their ability to capture subtle code details, consequently reducing the effectiveness of identifying similar sample pairs in LLM-generated outputs.
@@
Figure 1. An example illustrates how humans write correct code through thinking, and how models rely on a discriminator to select the correct code. The upper section illustrates the interaction with the code evaluation system, where the user attempts to submit the code twice. The lower section displays the correct code selection by a discriminator, where the model may not be able to generate the correct answer on its first attempt. Thus, it is necessary to employ a discriminator to enhance the reliability of the generated outputs.
@@
To keep the flexibility and enhance the detail-capturing capabilities of the code discriminator, we propose Condor, a non-execution-based discriminator that integrates general semantics with code details. To address the challenges that non-execution-based approaches face in capturing subtle code differences, we design an embedding-level contrastive learning strategy (Chen et al., 2020) and a data-level intermediate data mining strategy for Condor. In our embedding-level contrastive learning strategy, we first sample code pairs with high textual similarity for the same problem. Then, we design two loss functions during the training process to progressively bring the embeddings of correct code samples closer together, while increasing the embedding distance between correct and incorrect code snippets. Condor refines the base model’s code embeddings through contrastive learning to better distinguish between correct and incorrect code representations. For the data-level intermediate data mining strategy, we utilize data from multiple submissions for the same problem to identify intermediate data not covered by the existing training set. This strategy aims to capture intermediate data that the evaluation system may not have recorded during the user’s problem-solving process. For example, in the upper section of Figure 1, after the user’s thinking process (i.e., h4. Thinking), the user submits a new code snippet (i.e., submission 2) with modifications in two steps. After the modification in Step 1, the code still contains a bug. This partially modified version, which has not yet undergone the Step 2 modification, is referred to as partially fixed intermediate data. In real-world scenarios, users may submit a new code version only after careful consideration, and the partially fixed code for previous submissions may not always be explicitly recorded in the evaluation system. By leveraging these partially fixed buggy code snippets, Condor can incorporate the incremental details in the programmer’s thinking process to further enhance its discriminative capabilities in specific tasks. As shown in the lower section of Figure 1, by integrating these two strategies for capturing code details, Condor aims to automatically select the correct code from multiple outputs when the LLM struggles to generate the correct code on the first attempt.
@@
We propose two parameter-scale versions of Condor (i.e., 110M and 1.3B) and conduct extensive experiments to evaluate its effectiveness. We find that current datasets are insufficient for fully quantifying the discriminator’s ability to discern differences in code details (Chen et al., 2021b; Muennighoff et al., 2023). Thus, we construct the CodeNanoFix dataset, which includes a series of samples consisting of an algorithmic problem description, pairs of the code with subtle errors, and the corresponding correct version. First, we apply the Condor to the classification task to assess the model’s ability to recognize subtle code differences. The results show that general code generation models struggle to accurately distinguish these fine-grained differences, often resulting in imbalances in precision or recall, leading to suboptimal overall performance. In contrast, Condor excels across precision, recall, and F1-score, significantly outperforming other models. For example, Condor (110M) improves the F1-score from around 67% to 72.34%. Additionally, Condor (1.3B) demonstrates substantial improvements in discriminative capability when selecting the correct code from multiple generated outputs, achieving an increase in Pass@1 accuracy of Meta-Llama-3.1-Instruct (70B) from 52.64% to 62.63%. Second, to validate Condor’s generalization ability, we test it on the APPS (Hendrycks et al., 2021) and the MBPP (Austin et al., 2021) datasets. The results indicate that Condor consistently maintains strong performance across different datasets. For example, on the APPS dataset, Condor (1.3B) raises the Pass@1 of Meta-Llama-3.1 Instruct (70B) from 10.16% to 25.10%, and on MBPP, Condor (1.3B) improves Pass@1 from 71.40% to 75.20%. Finally, through ablation studies, we demonstrate that the two strategies for capturing code details in Condor are crucial to its overall performance, further validating their effectiveness in detecting subtle code differences.
@@
  • We propose Condor, a novel, non-execution-based code discriminator that incorporates embedding-level and data-level code details to discriminate and filter code generated by LLMs.
@@
2. Condor
2.1. Overview of Condor
Figure 2. The Condor overview consists of two main components: contrastive learning at the embedding level to capture code details (upper section), and data-level augmentation through intermediate code, which supplements code details that are not recorded in existing datasets (lower section).
@@ 2.1. Overview of Condor
2.2. Embedding-Level Contrast on Code Details
We employ a contrastive learning (Bui et al., 2021; Chen et al., 2024) strategy to make the model aware that textually similar yet functionally different code snippets should have distinct representations at the embedding level. Contrastive learning, known for enhancing the similarity between embeddings within the same category while reducing the similarity between those from different categories, is well-suited to our task. To fully utilize this mechanism, we design a similarity-based data sampling strategy tailored to code discrimination. By incorporating this design, we leverage contrastive learning to effectively capture subtle code differences, transforming them into meaningful parameter updates to refine the model’s code embedding representations.
2.2.1. Similarity-based Sampling
In contrastive learning, we focus on comparing samples with a high degree of textual similarity in the data sampling process, enabling the discriminator to differentiate between textually similar but functionally distinct code snippets. Each data sample in contrastive learning is represented as $\langle p,c_{a},c_{b},label_{a,b}\rangle$, where $c_{a}$ and $c_{b}$ are two textually similar code snippets, and $label_{a,b}$ indicates the relationship between them. Each contrastive learning sample takes into account both the problem description and the code representation, ensuring that the code embeddings are compared within the context of a specific problem. Based on whether the code can pass test cases (C: correct code that passes all test cases, E: error code that fails some test cases), the relationship between $c_{a}$ and $c_{b}$ can be classified into four categories: C-C, C-E, E-E, and E-C. Due to the symmetry between E-C and C-E in classification, we treat them as a single category. Since we primarily focus on the subtle differences between correct and incorrect code, we emphasize comparisons between C-C and E-C pairs. This helps the model recognize which code details affect functionality and which do not. Thus, the labels $label_{a,b}$ are divided into two key types: C-C (i.e., pairs with the same functionality) and E-C (pairs with different functionality). We assign a label of 1 to C-C pairs and 0 to E-C pairs and use these labeled data to train the supervised model for contrastive learning.
2.2.2. Contrastive Learning
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left[label_{a,b}\cdot d(x_{a}^{i},x_{b}^{i})^{2}+(1-label_{a,b})\cdot\max(0,m-d(x_{a}^{i},x_{b}^{i}))^{2}\right],$
where $N$ is the total number of training samples, $label_{a,b}$ is a binary label for each pair, and $label_{a,b}=1$ indicates the same functionality pairs (P-P) and $label_{a,b}=0$ indicates different functionality pairs (N-P). $d(x_{a}^{i},x_{b}^{i})$ is the Euclidean distance between the two embeddings $x_{a}^{i}$ and $x_{b}^{i}$. $m$ is the margin that ensures dissimilar pairs are separated by at least a certain distance.
@@
2.3. Data-Level Augmentation on Code Details
@@
2.3.1. Intermediate Code Details
@@ 2.3.2. Discriminating
2.4. CodeNanoFix Dataset
To better train a discriminator capable of detecting subtle code differences, it is essential to use data samples from real-world programming scenarios where programmers have written textually similar but functionally distinct code. Existing code generation datasets (Chen et al., 2021b; Austin et al., 2021; Hendrycks et al., 2021) are less suitable because these datasets lack real-world fine-grained modifications, and their training set relies on code generated by specific LLMs, making the discriminator’s effectiveness dependent on the particular LLM used. Additionally, identical functionality can be implemented in various ways, resulting in generated code that is not always textually similar. Therefore, we construct a new repair-related dataset, CodeNanoFix, comprising training, validation, and test sets to support the training of discriminators capable of detecting subtle changes in code.
2.4.1. Data Collection
CodeNanoFix consists of tuples of problem description, buggy code, and correct code, designed to evaluate human-written code with subtle differences. For example, Figure 1 illustrates two code snippets that are nearly identical in text but differ in functionality: Submission 1 on the left is incorrect (not passing the test cases), while submission 1 on the right is a correct solution corresponding to the problem description (passing the test cases). We collect the data samples from CodeNet (Puri et al., 2021), a large-scale collection of over 14 million code samples across over 50 programming languages, focusing on 4,000 problem-solving tasks from competitive programming. Each problem $p$ in the CodeNet dataset includes multiple submissions $S=(s_{1},s_{2},...)$ from different users and these submissions consist of both buggy and correct code snippets. Figure 1 depicts a user’s progression from an initial failed submission to a correctly revised version. We construct CodeNanoFix targeting examples where two submissions are textually similar but functionally different, specifically focusing on Python, the predominant language for evaluating code LLMs.
2.4.2. Data Cleaning
Data quality is crucial when building a code-related dataset. We observe several issues in the original CodeNet dataset that could potentially impact model training and evaluation. First, some code snippets in the dataset are mislabeled, such as code that fails certain test cases is incorrectly marked as functionally correct. To ensure label accuracy, we rerun all test cases and verify each label. We remove code samples where test results did not match the original labels, retaining only those where the test outputs are consistent with the original labels. Thus, we aim to maintain data consistency without introducing new labels, thereby avoiding potential label errors due to variations in test environments. Additionally, the problem descriptions in the original dataset are in different languages, which may interfere with the model’s understanding of the task. To address this, we use ChatGPT (OpenAI, 2024) to translate all problem descriptions into English for uniformity. These improvements enhance the dataset’s consistency and reliability, allowing it to reflect model performance in code discrimination tasks more accurately.
Table 1. The statistics of the CodeNanoFix dataset.

### v2

@@ Abstract
LLMs demonstrate significant potential across various software engineering tasks. However, they still face challenges in generating correct code on the first attempt when addressing complex requirements. Introducing a discriminator to select reliable outputs from multiple generated results is an effective way to enhance their reliability and stability. Currently, these discriminators fall into two categories: execution-based discriminators and non-execution-based discriminators. Execution-based discriminators face flexibility challenges due to difficulties in obtaining test cases and security concerns, while non-execution-based discriminators, although more flexible, struggle to capture subtle differences in code details. To maintain flexibility while improving the model’s ability to capture fine-grained code details, this paper proposes Condor. We first design contrastive learning to optimize the code representations of the base model, enabling it to reflect differences in code details. Then, we leverage intermediate data from the code modification process to further enrich the discriminator’s training data, enhancing its ability to discern code details. Experimental results indicate that on the subtle code difference dataset (i.e., CodeNanoFix), Condor significantly outperforms other discriminators in discriminative performance: Condor (1.3B) improves the discriminative F1 score of DeepSeek-Coder (1.3B) from 67% to 73%. In discriminating LLM-generated outputs, Condor (1.3B) and Condor (110M) raise the Pass@1 score of Llama-3.1-Instruct (70B) on the CodeNanoFix dataset from 52.64% to 62.63% and 59.64%, respectively. Moreover, Condor demonstrates strong generalization capabilities on the APPS, MBPP, and LiveCodeBench datasets. For example, Condor (1.3B) improves the Pass@1 of Llama-3.1-Instruct (70B) on the APPS dataset by 147.05%.
I Introduction
@@
With the widespread application of large language models (LLMs), their potential for tasks in software engineering (e.g., code generation and code repair) receives significant attention [guo2024deepseekcoder, dubey2024llama3, team2024codegemma, llm4understanding, chatgpt2024, scis_deeplearning]. However, they typically struggle to produce correct code on the first attempt when addressing complex requirements such as advanced algorithmic problems [apps, evaluating_code_quality, khan2023assessingpromisepitfallschatgpt, ji2023surveyhallucination]. Existing approaches typically employ an automated discriminator to select a correct result from multiple outputs generated by LLMs, improving the reliability and stability of LLMs when handling complex problems [coderanker].
@@
Currently, these approaches for code discrimination fall into two main categories: execution-based approaches and non-execution-based approaches. Execution-based approaches require running the code before making a discriminative decision, utilizing feedback from execution as additional detailed information to enhance the discriminator’s performance. However, execution-based models face challenges in flexibility. First, executing code poses challenges because (1) many real-world downstream reasoning tasks lack test cases [softwaretesting], (2) LLM-generated code may contain incomplete snippets [chen2021evaluatinglargelanguagemodels], and (3) the testing environments and other dependencies are not always accessible [coderanker, copilot]. Second, even when execution is possible, running LLM-generated code raises safety concerns. On the other hand, non-execution-based approaches do not execute code and rely solely on the details presented in the code text, providing greater flexibility and applicability across different application scenarios.
@@
However, existing non-execution-based discriminators typically leverage general code representation capabilities without adapting to the nuanced details of the programming process [coderanker, coder_reviewer_reranking]. Challenges remain in improving the discriminator’s effectiveness at capturing intricate code details. In the development scenario, the functionality of code often depends on key minor changes, such as variable updates, conditional logic modifications, or adjustments in function calls. For example, as shown in Figure 1, the upper section illustrates a problem description about finding a number K equally distant from two distinct numbers A and B. Submission 1 displays buggy code, while submission 2 shows the correct code after the user’s modification. Notably, only three characters differ between the two submissions (i.e., ‘(’, ‘)’, and ‘/’), yet the functionality of the first submission remains incorrect. Similarly, outputs generated by LLM can also contain errors caused by subtle differences [xu2024subtle_error, shi2024code_error], demonstrating that such mistakes are not unique to humans. For example, as shown below in Figure 1, the model (i.e., Llama-3.1-Instruct-70B) generates two outputs for Task 173 in the MBPP [mbpp] benchmark, one of which is incorrect due to a minor substitution, where ‘all’ is mistakenly replaced with ‘any’. Existing discriminators primarily rely on embeddings provided by other general-purpose code models as a basis for representation. This limits their ability to capture subtle code details, consequently reducing the effectiveness of identifying similar sample pairs in LLM-generated outputs.
@@
Figure 1: An example illustrates how humans write correct code through thinking, and how models rely on a discriminator to select the correct code. The upper section illustrates the interaction with the code evaluation system, where the user attempts to submit the code twice. The lower section displays the correct code selection by a discriminator, where the model may not be able to generate the correct answer on its first attempt. Thus, it is necessary to employ a discriminator to enhance the reliability of the generated outputs.
@@
To keep the flexibility and enhance the detail-capturing capabilities of the code discriminator, we propose Condor, a non-execution-based discriminator that integrates general semantics with code details. To address the challenges that non-execution-based approaches face in capturing subtle code differences, we design an embedding-level contrastive learning strategy [contrastive_learning] and a data-level intermediate data mining strategy for Condor. In our embedding-level contrastive learning strategy, we first sample code pairs with high textual similarity for the same problem. Then, we design two loss functions during the training process to progressively bring the embeddings of correct code samples closer together, while increasing the embedding distance between correct and incorrect code snippets. Condor refines the base model’s code embeddings through contrastive learning to better distinguish between correct and incorrect code representations. For the data-level intermediate data mining strategy, we utilize data from multiple submissions for the same problem to identify intermediate data not covered by the existing training set. This strategy aims to capture intermediate data that the evaluation system may not have recorded during the user’s problem-solving process. For example, in the upper section of Figure 1, after the user’s thinking process (i.e., h4. Thinking), the user submits a new code snippet (i.e., submission 2) with modifications in two steps. After the modification in Step 1, the code still contains a bug. This partially modified version, which has not yet undergone the Step 2 modification, is referred to as partially fixed intermediate data. In real-world scenarios, users may submit a new code version only after careful consideration, and the partially fixed code for previous submissions may not always be explicitly recorded in the evaluation system. By leveraging these partially fixed buggy code snippets, Condor can incorporate the incremental details in the programmer’s thinking process to further enhance its discriminative capabilities in specific tasks. As shown in the lower section of Figure 1, by integrating these two strategies for capturing code details, Condor aims to automatically select the correct code from multiple outputs when the LLM struggles to generate the correct code on the first attempt.
@@
We implement two Condor versions (i.e., 110M and 1.3B) to balance performance and practicality: the smaller model enables low-latency ranking in resource-constrained settings [codebert2020, coderanker, codet52021], while the larger model offers stronger accuracy, providing a better trade-off than single-shot inference with much larger models [guo2024deepseekcoder, grammarcoder]. When training a discriminator, we find that current datasets are insufficient for fully quantifying the discriminator’s ability to discern differences in code details [humaneval, commitpack_muennighoff2023octopack]. Thus, we construct the CodeNanoFix dataset, which includes a series of samples consisting of an algorithmic problem description, pairs of code with subtle errors, and the corresponding correct version. First, we apply the Condor to the classification task, evaluating its ability to recognize subtle code differences. The results show that general code generation models struggle to accurately distinguish these fine-grained differences, often resulting in imbalances in precision or recall, leading to suboptimal overall performance. In contrast, Condor excels across precision, recall, and F1-score, significantly outperforming other models. For example, Condor (110M) improves the F1-score from around 67% to 72.34%. Additionally, Condor (1.3B) demonstrates substantial improvements in discriminative capability when selecting the correct code from multiple generated outputs, achieving an increase in Pass@1 accuracy of Llama-3.1-Instruct (70B) from 52.64% to 62.63%. Second, to validate Condor’s generalization ability, we test it on the APPS [apps], the MBPP [mbpp], and the LiveCodeBench [jain2024livecodebench] datasets. The results indicate that Condor consistently maintains strong performance across different datasets. For example, on the APPS dataset, Condor (1.3B) raises the Pass@1 of Llama-3.1 Instruct (70B) from 10.16% to 25.10%, on MBPP, Condor (1.3B) improves Pass@1 from 71.40% to 75.20%, and on LiveCodeBench, Condor (1.3B) improves Pass@1 from 28.97% to 30.90%. Finally, through ablation studies, we demonstrate that the two strategies for capturing code details in Condor are crucial to its overall performance, further validating their effectiveness in detecting subtle code differences.
@@
  • We propose Condor [condor_github], a novel, non-execution-based code discriminator that incorporates embedding-level and data-level code details to discriminate and filter code generated by LLMs.
@@
II Condor
II-A Overview of Condor
Figure 2: The Condor overview consists of two main components: contrastive learning at the embedding level to capture code details (upper section), and data-level augmentation through intermediate code, which supplements code details that are not recorded in existing datasets (lower section). The ‘C’ denotes the correct code that passes all test cases, while ‘E’ indicates the error code that fails some test cases.
@@ II-A Overview of Condor
II-B Embedding-Level Contrast on Code Details
We employ a contrastive learning [contrastive_1, contrastive_2, zhao2024spotting, chen2024improving] strategy to make the model aware that textually similar yet functionally different code snippets should have distinct representations at the embedding level. Contrastive learning, known for enhancing the similarity between embeddings within the same category while reducing the similarity between those from different categories, is well-suited to our task. To fully utilize this mechanism, we design a similarity-based data sampling strategy tailored to code discrimination. By incorporating this design, we leverage contrastive learning to effectively capture subtle code differences, transforming them into meaningful parameter updates to refine the model’s code embedding representations.
II-B1 Similarity-based Sampling
In contrastive learning, we focus on comparing samples with a high degree of textual similarity in the data sampling process, enabling the discriminator to differentiate between textually similar but functionally distinct code snippets. Each data sample in contrastive learning is represented as $\langle p,c_{a},c_{b},label_{a,b}\rangle$, where $p$ represent the problem description, $c_{a}$ and $c_{b}$ are two textually similar code snippets, and $label_{a,b}$ indicates the relationship between them. Each contrastive learning sample takes into account both the problem description and the code representation, ensuring that the code embeddings are compared within the context of a specific problem. Based on whether the code can pass test cases (C: correct code that passes all test cases, E: error code that fails some test cases), the relationship between $c_{a}$ and $c_{b}$ can be classified into four categories: C-C, C-E, E-E, and E-C. Due to the symmetry between E-C and C-E in classification, we treat them as a single category. Since we primarily focus on the subtle differences between correct and incorrect code, we emphasize comparisons between C-C and E-C pairs. This helps the model recognize which code details affect functionality and which do not. Thus, the labels $label_{a,b}$ are divided into two key types: C-C (i.e., pairs with the same functionality) and E-C (pairs with different functionality). We assign a label of 1 to C-C pairs and 0 to E-C pairs and use these labeled data to train the supervised model for contrastive learning.
II-B2 Contrastive Learning
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\bigg[label_{a,b}\cdot d(x_{a}^{i},x_{b}^{i})^{2}\\ +(1-label_{a,b})\cdot\max(0,m-d(x_{a}^{i},x_{b}^{i}))^{2}\bigg],$
where $N$ is the total number of training samples, $label_{a,b}$ is a binary label for each pair, and $label_{a,b}=1$ indicates the same functionality pairs (C-C) and $label_{a,b}=0$ indicates different functionality pairs (E-C). $d(x_{a}^{i},x_{b}^{i})$ is the Euclidean distance between the two embeddings $x_{a}^{i}$ and $x_{b}^{i}$. $m$ is the margin that ensures dissimilar pairs are separated by at least a certain distance.
@@
II-C Data-Level Augmentation on Code Details
@@
II-C1 Intermediate Code Details
@@ II-C2 Discriminating
II-D CodeNanoFix Dataset
To better train a discriminator capable of detecting subtle code differences, it is essential to use data samples from real-world programming scenarios where programmers have written textually similar but functionally distinct code. Existing code generation datasets [humaneval, mbpp, apps] are less suitable because these datasets lack real-world fine-grained modifications, and their training set relies on code generated by specific LLMs, making the discriminator’s effectiveness dependent on the particular LLM used. Additionally, identical functionality can be implemented in various ways, resulting in generated code that is not always textually similar. Therefore, we construct a new repair-related dataset, CodeNanoFix, comprising training, validation, and test sets to support the training of discriminators capable of detecting subtle changes in code.
II-D1 Data Collection
[…]
II-D2 Data Cleaning
Data quality is crucial when building a code-related dataset. We observe several issues in the original CodeNet dataset that could potentially impact model training and evaluation. First, some code snippets in the dataset are mislabeled, such as code that fails certain test cases is incorrectly marked as functionally correct. To ensure label accuracy, we rerun all test cases and verify each label. We remove code samples where test results did not match the original labels, retaining only those where the test outputs are consistent with the original labels. Thus, we aim to maintain data consistency without introducing new labels, thereby avoiding potential label errors due to variations in test environments. Additionally, the problem descriptions in the original dataset are in different languages, which may interfere with the model’s understanding of the task. To address this, we use ChatGPT [chatgpt2024] to translate all problem descriptions into English for uniformity. These improvements enhance the dataset’s consistency and reliability, allowing it to reflect model performance in code discrimination tasks more accurately.
TABLE I: The statistics of the CodeNanoFix dataset.